OpenStudy (ny,ny):

a farmer grows only sunflowers and flax on his 240 acre farm. this year he wants to plant 80 more acres of sunflowers than of flax. how many acres of each crop does the farmer need to plant?

OpenStudy (ny,ny):

sorry, i left some of the directions out. define variables and write a system of equations for each situation. solve using substitution.

OpenStudy (ny,ny):

please help, i really need to get to bed. D:

OpenStudy (raden):

x+y=240 ... (1) x=y+80 ...(2) subst (2) to (1)

OpenStudy (ny,ny):

x= flax y= sunflower?

OpenStudy (raden):

no, x=sunflower y= flax

OpenStudy (ny,ny):

ohh. ok. i want to use s and f. much easier to determine between them. f=80?

OpenStudy (raden):

ok, s+f=240 and s=f+80 so, f+80+f=240 2f+80=240 2f=240-80 2f=160 f=160/2=80 u are correct

OpenStudy (ny,ny):

s=160?

OpenStudy (raden):

just put f=80 to s=f+80=80+80=160 correct again :)

OpenStudy (ny,ny):

thank you! you are awesome :D can you please please please help me with another question? i closed it already though because there was no replies on it.

OpenStudy (raden):

welcome... of course, post ur question